# Learning multi-modal generative models with permutation-invariant encoders and tighter variational bounds

### v2

@@ 4.1 Identifiability
Identifiability of parameters and latent variables in latent structure models is a classic problem (Koopmans and Reiersol, 1950; Kruskal, 1976; Allman et al., 2009), that has been studied increasingly for non-linear latent variable models, e.g., for ICA (Hyvarinen and Morioka, 2016; Hälvä and Hyvarinen, 2020; Hälvä et al., 2021), VAEs (Khemakhem et al., 2020a; Zhou and Wei, 2020; Wang et al., 2021; Moran et al., 2021; Lu et al., 2022; Kim et al., 2023), EBMs (Khemakhem et al., 2020b), flow-based (Sorrenson et al., 2020) or mixture models (Kivva et al., 2022).
Non-linear generative models are generally unidentifiable without imposing some structure (Hyvärinen and Pajunen, 1999; Xi and Bloem-Reddy, 2022). Yet, identifiability up to some ambiguity can be achieved in some conditional models based on observed auxiliary variables and injective decoder functions wherein the prior density is conditional on auxiliary variables. Observations from different modalities can act as auxiliary variables to obtain identifiability of conditional distributions given some modality subset under analogous assumptions.
@@
In the iVAE model (Khemakhem et al., 2020a), the latent variable distribution $p_{\theta}(z|x_{1})$ is independently modulated via an auxiliary variable $X_{1}=U$. Instead of interpreting this distribution as a (conditional) prior density, we view it as a posterior density given the first modality $X_{1}$. Khemakhem et al. (2020a) estimate a model for another modality $X_{2}$ by lower bounding $\log p_{\theta}(x_{2}|x_{1})$ via $\mathcal{L}_{\setminus\left\{1\right\}}$ under the assumption that $q_{\phi}(z|x_{1})$ is given by the prior density $p_{\theta}(z|x_{1})$. Similarly, Mita et al. (2021) optimise $\log p_{\theta}(x_{1},x_{2})$ by a double VAE bound that reduces to $\mathcal{L}$ for a masking distribution $\rho(s_{1},s_{2})=(\delta_{1}\otimes\delta_{0})(s_{1},s_{2})$ that always masks the modality $X_{2}$ and choosing to parameterize separate encoding functions for different conditioning sets. Our bound thus generalizes these procedures to multiple modalities in a scalable way.
We are interested in identifiability, conditional on having observed some non-empty modality subset $\mathcal{S}\subset\mathcal{M}$. For illustration, we translate an identifiability result from the uni-modal iVAE setting in Lu et al. (2022), which does not require the conditional independence assumption from Khemakhem et al. (2020a). We assume that the encoding distribution $q_{\phi}(z|x_{\mathcal{S}})$ approximates the true posterior $p_{\theta}(z|x_{\mathcal{S}})$ and belongs to a strongly exponential family, i.e.,
@@
This result follows from Theorem 4 in Lu et al. (2022). Note that $p_{\theta_{\mathcal{S}}}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})=p_{\tilde{\theta}_{\mathcal{S}}}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ for all $x\in\mathsf{X}$ implies with the regularity assumption on $\varphi_{\setminus\mathcal{S},\epsilon}$ that the transformed variables $Z=f_{\setminus\mathcal{S}}^{-1}(X_{\setminus\mathcal{S}})$ and $\tilde{Z}=\tilde{f}_{\setminus\mathcal{S}}^{-1}(X_{\setminus\mathcal{S}})$ have the same density function conditional on $X_{\mathcal{S}}$.
@@
Note that (13) and (12) imply that each modality has a Lebesgue density under the generative model. This assumption may not hold for some modalities, such as imaging data that can be supported (closely) on a lower-dimensional manifold (Roweis and Saul, 2000), causing issues in likelihood-based methods such as VAEs (Dai and Wipf, 2018; Loaiza-Ganem et al., 2022). Moreover, different conditioning sets or modalities may result in different dimensions of the underlying manifold for conditional data (Zheng et al., 2022). Some two-step approaches (Dai and Wipf, 2018; Zheng et al., 2022) first estimate the dimension $r$ of the ground-truth manifold as a function of the encoder variance relative to the variance under the (conditional) prior for each latent dimension $i$, $i\in[D]$, with $r\leq D$. It would, therefore, be interesting to analyze in future work if more flexible aggregation schemes that do not impose strong biases on the variance components of the encoder can better learn the manifold dimensions in conditional or multi-modal models following an analogous two-step approach.
Recall that the identifiability considered here concerns parameters of the multi-modal posterior distribution. We believe that our inference approach is beneficial for this type of identifiability because (a) unlike some other variational bounds, the posterior is the optimal variational distribution with $\mathcal{L}_{\setminus\mathcal{S}}(x)$ being an approximate lower bound on $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$, and (b) the trainable aggregation schemes can be more flexible for approximating the optimal encoding distribution.
@@ 4.2 Mixture models
An alternative to the choice of uni-modal prior densities $p_{\theta}$ has been to use Gaussian mixture priors (Johnson et al., 2016; Jiang et al., 2017; Dilokthanakul et al., 2016) or more flexible mixture models (Falck et al., 2021). Following previous work, we include a latent cluster indicator variable $c\in[K]$ that indicates the mixture component out of $K$ possible mixtures with augmented prior $p_{\theta}(c,z)=p_{\theta}(c)p_{\theta}(z|c)$. The classic example is $p_{\theta}(c)$ being a categorical distribution and $p_{\theta}(z|c)$ a Gaussian with mean $\mu_{c}$ and covariance matrix $\Sigma_{c}$. Similar to Falck et al. (2021) that use an optimal variational factor in a mean-field model, we use an optimal factor of the cluster indicator in a structured variational density $q_{\phi}(c,z|x_{\mathcal{S}})=q_{\phi}(z|x_{\mathcal{S}})q_{\phi}(c|z,x_{\mathcal{S}})$ with $q_{\phi}(c|z,x_{\mathcal{S}})=p_{\theta}(c|z)$. Appendix G details how one can optimize an augmented multi-modal bound. Concurrent work (Palumbo et al., 2024) considered a similar optimal variational factor for a discrete mixture model under a MoE aggregation.
@@ 4.3 Missing modalities
In practical applications, modalities can be missing for different data points. We describe this missingness pattern by missingness mask variables $m_{s}\in\{0,1\}$ where $m_{s}=1$ indicates that observe modality $s$, while $m_{s}=0$ means it is missing. The joint generative model that extends (17) will be of the form $p_{\theta}(z,x,m)=p_{\theta}(z)\prod_{s\in\mathcal{M}}p_{\theta}(x_{s}|z)p_{\theta}(m|x)$ for some distribution $p_{\theta}(m|x)$ over the mask variables $m=(m_{s})_{s\in\mathcal{M}}$. For $\mathcal{S}\subset\mathcal{M}$, we denote by $x_{\mathcal{S}}^{o}=\{x_{s}\colon m_{s}=1,s\in\mathcal{S}\}$ and $x_{\mathcal{S}}^{m}=\{x_{s}\colon m_{s}=0,s\in\mathcal{S}\}$ the set of observed, respectively missing, modalities. The full likelihood of the observed and missingness masks becomes then $p_{\theta}(x_{\mathcal{S}}^{o},m)=\int p_{\theta}(z)\prod_{s\in\mathcal{S}}p_{\theta}(x_{s}|z)p_{\theta}(m|x)\mathrm{d}x_{s}^{m}\mathrm{d}z$. If $p_{\theta}(m|x)$ does not depend on the observations, that is, observations are missing completely at random (Rubin, 1976), then the missingness mechanisms $p_{\theta}(m|x)$ for inference approaches maximizing $p_{\theta}(x^{o},m)$ can be ignored. Consequently, one can instead concentrate on maximizing $\log p_{\theta}(x^{o})$ only, based on the joint generative model $p_{\theta}(z,x^{o})=p_{\theta}(z)\prod_{\left\{s\in\mathcal{M}\colon m_{s}=1\right\}}p_{\theta}(x_{s}|z)$. In particular, one can employ the variational bounds above by considering only the observed modalities. Since masking operations are readily supported for the considered permutation-invariant models, appropriate imputation strategies (Nazabal et al., 2020; Ma et al., 2019) for the encoded features of the missing modalities are not necessarily required. Settings allowing for not (completely) at random missingness have been considered in the uni-modal case, for instance, in Ipsen et al. (2021); Ghalebikesabi et al. (2021); Gong et al. (2021), and we leave multi-modal extensions thereof for future work for a given aggregation approach.
@@ 5 Experiments
We present a series of numerical experiments that illustrate the benefits of learning more flexible aggregation models and that optimizing our variational bound - called masked in the subsequent Figures - leads to higher log-likelihood values. Moreover, we demonstrate that our variational bound results in models that differ in their information theoretic quantities compared to models trained with a mixture-based bound. Recall that the full reconstruction log-likelihood is the negative full distortion $-D_{\mathcal{M}}$ based on all modalities, while the full rate $R_{\mathcal{M}}$ is the averaged KL between the encoding distribution of all modalities and the prior. Across different datasets, we observe that our bound leads to higher reconstructions $-D_{\mathcal{M}}$ for and given all modalities with increased full rates $R_{\mathcal{M}}$. Note that mixture-based bounds optimize directly for the cross-reconstruction log-likelihood $D^{c}_{\setminus\mathcal{S}}$, see Remark 9, and do not contain a cross-rate term $R_{\setminus\mathcal{S}}$, i.e. the KL between the encoding distribution for all modalities relative to a modality subset, as a regulariser, in contrast to our bound (Lemma 7 and Corollary 8). This observation is confirmed empirically by higher cross reconstructions with increased cross rates for mixture-based bounds.
@@ 5.1 Linear multi-modal VAEs
The relationship between uni-modal VAEs and probabilistic PCA (Tipping and Bishop, 1999) has been studied in previous work (Dai et al., 2018; Lucas et al., 2019; Rolinek et al., 2019; Huang et al., 2020; Mathieu et al., 2019). We analyze how different multi-modal fusion schemes and multi-modal variational bounds affect (a) the learned generative model in terms of its true marginal log-likelihood (LLH) and (b) the latent representations in terms of information-theoretic quantities and identifiability. To evaluate the (weak) identifiability of the method, we follow Khemakhem et al. (2020a, b) to compute the mean correlation coefficient (MCC) between the true latent variables $Z$ and samples from the variational distribution $q_{\phi}(\cdot|x_{\mathcal{M}})$ after an affine transformation using CCA.
@@
for a standard Gaussian prior $Z\sim\mathcal{N}(0,\operatorname{I})$ and independent noise variables $U_{1},U_{2}\sim\mathcal{N}(0,1)$. Note that the second modality is more noisy compared to the first one. The results in Table 1 for the obtained log-likelihood values show first that learnable aggregation models yield higher log-likelihoods [^7] We found that a PoE model can have numerical issues here., and second that our bound yields higher log-likelihood values compared to mixture-based bounds for any given fixed aggregation model. We also compute various information theoretic quantities, confirming that our bound leads to higher full reconstructions at higher full rates and lower cross reconstructions at lower cross rates, compared to mixture-based bounds. More flexible aggregation schemes increase the full and cross reconstruction for any given bound, while not necessarily increasing the full or cross rates, i.e., they can result in an improved point within a rate-distortion curve for some configurations.
@@
Relative LLH gap Full Reconstruction Full Rates Cross Reconstruction Cross Rates
Aggregation ours mixture ours mixture ours mixture ours mixture ours mixture
@@
We consider $M=5$ modalities following multi-variate Gaussian laws. We consider generative models wherin all latent variables are shared across all modalities, as well as generative models where only parts of the latent variables are shared across all modalities, while the remaining latent variables being modality specific. The setting of private latent variables can be incorporated by imposing sparsity structures on the decoding matrices and allows us to analyze scenarios with considerable modality-specific variation described through private latent variables. We provide more details about the data generation mechanisms in Appendix J. For illustration, we use multi-modal encoders with shared latent variables using invariant aggregations in the first case, and multi-modal encoders that utilise additionally equivariant aggregations for the private latent variables in the second case. Results in Table 2 suggest that more flexible aggregation schemes improve the LLH and the identifiability for both variational objectives. Furthermore, our new bound yields higher LLH for a given aggregation scheme.
Table 2: Gaussian model with five modalities: Relative difference of true LLH to the learned LLH. MCC to true latent. The generative model for the invariant aggregation schemes are using dense decoders, whereas the ground truth model for the permutation-equivariant encoders use sparse decoders to account for private latent variables.
@@
Our bound Mixture bound Our bound Mixture bound

### v3

@@ 4.1 Identifiability
Identifiability of parameters and latent variables in latent structure models is a classic problem [Koopmans and Reiersol, 1950, Kruskal, 1976, Allman et al., 2009], that has been studied increasingly for non-linear latent variable models, e.g., for ICA [Hyvarinen and Morioka, 2016, Hälvä and Hyvarinen, 2020, Hälvä et al., 2021], VAEs [Khemakhem et al., 2020a, Zhou and Wei, 2020, Wang et al., 2021, Moran et al., 2021, Lu et al., 2022, Kim et al., 2023], EBMs [Khemakhem et al., 2020b], flow-based [Sorrenson et al., 2020] or mixture models [Kivva et al., 2022].
Non-linear generative models are generally unidentifiable without imposing some structure [Hyvärinen and Pajunen, 1999, Xi and Bloem-Reddy, 2022]. Yet, identifiability up to some ambiguity can be achieved in some conditional models based on observed auxiliary variables and injective decoder functions wherein the prior density is conditional on auxiliary variables. Observations from different modalities can act as auxiliary variables to obtain identifiability of conditional distributions given some modality subset under analogous assumptions.
@@
In the iVAE model [Khemakhem et al., 2020a], the latent variable distribution $p_{\theta}(z|x_{1})$ is independently modulated via an auxiliary variable $X_{1}=U$. Instead of interpreting this distribution as a (conditional) prior density, we view it as a posterior density given the first modality $X_{1}$. Khemakhem et al. [2020a] estimate a model for another modality $X_{2}$ by lower bounding $\log p_{\theta}(x_{2}|x_{1})$ via $\mathcal{L}_{\setminus\left\{1\right\}}$ under the assumption that $q_{\phi}(z|x_{1})$ is given by the prior density $p_{\theta}(z|x_{1})$. Similarly, Mita et al. [2021] optimize $\log p_{\theta}(x_{1},x_{2})$ by a double VAE bound that reduces to $\mathcal{L}$ for a masking distribution $\rho(s_{1},s_{2})=(\delta_{1}\otimes\delta_{0})(s_{1},s_{2})$ that always masks the modality $X_{2}$ and choosing to parameterize separate encoding functions for different conditioning sets. Our bound thus generalizes these procedures to multiple modalities in a scalable way.
We are interested in identifiability, conditional on having observed some non-empty modality subset $\mathcal{S}\subset\mathcal{M}$. For illustration, we translate an identifiability result from the uni-modal iVAE setting in Lu et al. [2022], which does not require the conditional independence assumption from Khemakhem et al. [2020a]. We assume that the encoding distribution $q_{\phi}(z|x_{\mathcal{S}})$ approximates the true posterior $p_{\theta}(z|x_{\mathcal{S}})$ and belongs to a strongly exponential family, i.e.,
@@
This result follows from Theorem 4 in Lu et al. [2022]. Note that $p_{\theta_{\mathcal{S}}}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})=p_{\tilde{\theta}_{\mathcal{S}}}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$ for all $x\in\mathsf{X}$ implies with the regularity assumption on $\varphi_{\setminus\mathcal{S},\epsilon}$ that the transformed variables $Z=f_{\setminus\mathcal{S}}^{-1}(X_{\setminus\mathcal{S}})$ and $\tilde{Z}=\tilde{f}_{\setminus\mathcal{S}}^{-1}(X_{\setminus\mathcal{S}})$ have the same density function conditional on $X_{\mathcal{S}}$.
@@
Note that (14) and (13) imply that each modality has a Lebesgue density under the generative model. This assumption may not hold for some modalities, such as imaging data that can be supported (closely) on a lower-dimensional manifold [Roweis and Saul, 2000], causing issues in likelihood-based methods such as VAEs [Dai and Wipf, 2018, Loaiza-Ganem et al., 2022]. Moreover, different conditioning sets or modalities may result in different dimensions of the underlying manifold for conditional data [Zheng et al., 2022]. Some two-step approaches [Dai and Wipf, 2018, Zheng et al., 2022] first estimate the dimension $r$ of the ground-truth manifold as a function of the encoder variance relative to the variance under the (conditional) prior for each latent dimension $i$, $i\in[D]$, with $r\leq D$. It would, therefore, be interesting to analyze in future work if more flexible aggregation schemes that do not impose strong biases on the variance components of the encoder can better learn the manifold dimensions in conditional or multi-modal models following an analogous two-step approach.
Recall that the identifiability considered here concerns parameters of the multi-modal posterior distribution and the conditional generative distribution. It is thus preliminary to estimation and only concerns the generative model and not the inference approach. However, both the multi-modal posterior distribution and the conditional generative distribution are intractable. In practice, we thus replace them with approximations. We believe that our inference approach is beneficial for this type of identifiability when making these variational approximations because (a) unlike some other variational bounds, the posterior is the optimal variational distribution with $\mathcal{L}_{\setminus\mathcal{S}}(x)$ being an approximation of a lower bound on $\log p_{\theta}(x_{\setminus\mathcal{S}}|x_{\mathcal{S}})$, see Remark 10, and (b) the trainable aggregation schemes can be more flexible for approximating the optimal encoding distribution.
@@ 4.2 Mixture models
An alternative to the choice of uni-modal prior densities $p_{\theta}$ has been to use Gaussian mixture priors [Johnson et al., 2016, Jiang et al., 2017, Dilokthanakul et al., 2016] or more flexible mixture models [Falck et al., 2021]. Following previous work, we include a latent cluster indicator variable $c\in[K]$ that indicates the mixture component out of $K$ possible mixtures with augmented prior $p_{\theta}(c,z)=p_{\theta}(c)p_{\theta}(z|c)$. The classic example is $p_{\theta}(c)$ being a categorical distribution and $p_{\theta}(z|c)$ a Gaussian with mean $\mu_{c}$ and covariance matrix $\Sigma_{c}$. Similar to Falck et al. [2021] that use an optimal variational factor in a mean-field model, we use an optimal factor of the cluster indicator in a structured variational density $q_{\phi}(c,z|x_{\mathcal{S}})=q_{\phi}(z|x_{\mathcal{S}})q_{\phi}(c|z,x_{\mathcal{S}})$ with $q_{\phi}(c|z,x_{\mathcal{S}})=p_{\theta}(c|z)$. Appendix G details how one can optimize an augmented multi-modal bound. Concurrent work [Palumbo et al., 2024] considered a similar optimal variational factor for a discrete mixture model under a MoE aggregation.
@@ 4.3 Missing modalities
In practical applications, modalities can be missing for different data points. We describe this missingness pattern by missingness mask variables $m_{s}\in\{0,1\}$ where $m_{s}=1$ indicates that observe modality $s$, while $m_{s}=0$ means it is missing. The joint generative model with missing modalities will be of the form $p_{\theta}(z,x,m)=p_{\theta}(z)\prod_{s\in\mathcal{M}}p_{\theta}(x_{s}|z)p_{\theta}(m|x)$ for some distribution $p_{\theta}(m|x)$ over the mask variables $m=(m_{s})_{s\in\mathcal{M}}$. For $\mathcal{S}\subset\mathcal{M}$, we denote by $x_{\mathcal{S}}^{o}=\{x_{s}\colon m_{s}=1,s\in\mathcal{S}\}$ and $x_{\mathcal{S}}^{m}=\{x_{s}\colon m_{s}=0,s\in\mathcal{S}\}$ the set of observed, respectively missing, modalities. The full likelihood of the observed and missingness masks becomes then $p_{\theta}(x_{\mathcal{S}}^{o},m)=\int p_{\theta}(z)\prod_{s\in\mathcal{S}}p_{\theta}(x_{s}|z)p_{\theta}(m|x)\mathrm{d}x_{s}^{m}\mathrm{d}z$. If $p_{\theta}(m|x)$ does not depend on the observations, that is, observations are missing completely at random [Rubin, 1976], then the missingness mechanisms $p_{\theta}(m|x)$ for inference approaches maximizing $p_{\theta}(x^{o},m)$ can be ignored. Consequently, one can instead concentrate on maximizing $\log p_{\theta}(x^{o})$ only, based on the joint generative model $p_{\theta}(z,x^{o})=p_{\theta}(z)\prod_{\left\{s\in\mathcal{M}\colon m_{s}=1\right\}}p_{\theta}(x_{s}|z)$. In particular, one can employ the variational objectives above by considering only the observed modalities. Since masking operations are readily supported for the considered permutation-invariant models, appropriate imputation strategies [Nazabal et al., 2020, Ma et al., 2019] for the encoded features of the missing modalities are not necessarily required. Settings allowing for not (completely) at random missingness have been considered in the uni-modal case, for instance, in Ipsen et al. [2021], Ghalebikesabi et al. [2021], Gong et al. [2021], and we leave multi-modal extensions thereof for future work for a given aggregation approach.
@@ 5 Experiments
We conduct a series of numerical experiments to illustrate the effects of different variational objectives and aggregation schemes. Recall that the full reconstruction log-likelihood is the negative full distortion $-D_{\mathcal{M}}$ based on all modalities, while the full rate $R_{\mathcal{M}}$ is the averaged KL between the encoding distribution of all modalities and the prior. Note that mixture-based bounds maximize directly the cross-modal log-likelihood $-D^{c}_{\setminus\mathcal{S}}$, see (4), and do not contain a cross-rate term $R_{\setminus\mathcal{S}}$, i.e. the KL between the encoding distribution for all modalities relative to a modality subset, as a regulariser, in contrast to our objective (Lemma 7 and Corollary 8). The log-likelihood should be higher if a generative model is able to capture modality-specific information for models trained with $\beta=1$. For arbitrary $\beta$, we can take a rate-distortion perspective and look at how different generative models self-reconstruct all modalities, i.e., the full reconstruction term $-D_{\mathcal{M}}$, relative to the KL-divergence between the multi-modal encoding distribution and the prior, i.e. $R_{\mathcal{M}}$. This corresponds to a rate-distortion analysis of a VAE that merges all modalities into a single modality. A high full-reconstruction term is thus indicative of the encoder and decoder being able to reconstruct all modalities precisely so that they do not produce an average prediction. Note that neither our objective nor the mixture-based bound optimize for the full-reconstruction term directly.
@@ 5.1 Linear multi-modal VAEs
The relationship between uni-modal VAEs and probabilistic PCA [Tipping and Bishop, 1999] has been studied in previous work [Dai et al., 2018, Lucas et al., 2019, Rolinek et al., 2019, Huang et al., 2020, Mathieu et al., 2019]. We analyze how different multi-modal fusion schemes and multi-modal variational objectives affect (a) the learned generative model in terms of its true marginal log-likelihood (LLH) and (b) the latent representations in terms of information-theoretic quantities and identifiability. To evaluate the (weak) identifiability of the method, we follow Khemakhem et al. [2020a, b] to compute the mean correlation coefficient (MCC) between the true latent variables $Z$ and samples from the variational distribution $q_{\phi}(\cdot|x_{\mathcal{M}})$ after an affine transformation using CCA.
@@
for a standard Gaussian prior $Z\sim\mathcal{N}(0,\operatorname{I})$ and independent noise variables $U_{1},U_{2}\sim\mathcal{N}(0,1)$. Note that the second modality is more noisy compared to the first one. The results in Table 1 for the obtained log-likelihood values show first that learnable aggregation models yield higher log-likelihoods [^6] We found that a PoE model can have numerical issues here., and second that our bound yields higher log-likelihood values compared to mixture-based bounds for any given fixed aggregation model. We also compute various information theoretic quantities, confirming that our bound leads to higher full reconstructions at higher full rates and lower cross predictions at lower cross rates compared to mixture-based bounds. More flexible aggregation schemes increase the full and cross predictions for any given bound while not necessarily increasing the full or cross rates, i.e., they can result in an improved point within a rate-distortion curve for some configurations.
@@
Relative LLH gap Full Reconstruction Full Rates Cross Prediction Cross Rates
Aggregation our obj. mixture bound our obj. mixture bound our obj. mixture bound our ob. mixture bound our obj. mixture bound
@@
We consider $M=5$ modalities following multi-variate Gaussian laws. We consider generative models, where all latent variables are shared across all modalities, as well as generative models, where only parts of the latent variables are shared across all modalities, while the remaining latent variables are modality-specific. The setting of private latent variables can be incorporated by imposing sparsity structures on the decoding matrices and allows us to analyze scenarios with considerable modality-specific variation described through private latent variables. We provide more details about the data generation mechanisms in Appendix J. For illustration, we use multi-modal encoders with shared latent variables using invariant aggregations in the first case and multi-modal encoders that utilize additional equivariant aggregations for the private latent variables in the second case. Results in Table 2 suggest that more flexible aggregation schemes improve the LLH and the identifiability for both variational objectives. Furthermore, our new bound yields higher LLH for a given aggregation scheme.
Table 2: Gaussian model with five modalities: Relative difference of true LLH to the learned LLH. MCC to true latent. The generative model for the invariant aggregation schemes uses dense decoders, whereas the ground truth model for the permutation-equivariant encoders uses sparse decoders to account for private latent variables. We report mean values with standard deviations in parentheses over five independent runs.
@@
Proposed objective Mixture bound Proposed objective Mixture bound